OpenStudy (psycho45):

aider s’il vous plait 4+1/3x=5/12x−6

OpenStudy (psycho45):

NVM, que je l’ai eu

OpenStudy (welshfella):

Bon!

OpenStudy (psycho45):

u peut m’aider avec celui-ci \[-\frac{ 2 }{ 5 }x-\frac{ 3 }{ 4 }x=-\frac{ 99 }{ 10 }+\frac{ 1 }{ 2 }x\]